Please help me solve this equation. {How many grams of 4% salt solution must be mixed with 80 grams of 11% salt solution to make an 8% salt solution?}
Answer by stanbon(75887)

Please help me solve this equation. {How many grams of 4% salt solution must be mixed with 80 grams of 11% salt solution to make an 8% salt solution?}
-------------
Let amount of 4% solution be "x" ; amount of salt in it is 0.04x grams
Amount of 11% solution is 80 grams; amount of salt in it is 0.11(80)=8.8 grms.
-----------
EQUATIOn:
salt + salt = 0.08(80+x)
0.04x+8.8=6.4+0.08x
-0.04x=-2.4
x=60 (amount of 4% solution that must be mixed)
Cheers,
Stan H.